Find the probability that the number of boxes delivered will be atleast 37 but not more than 50
1
Expert's answer
2021-09-23T00:08:48-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Day & Number\ of\\
 & boxes \ (X)\\ \hline
 1 & 35 \\
 \hdashline
 2 & 37 \\
 \hdashline
 3 & 50 \\
 \hdashline
 4 & 45 \\
 \hdashline
 5 & 37 \\
 \hdashline
 6 & 45 \\
 \hdashline
 7 & 40 \\
 \hdashline
 8 & 42 \\
 \hdashline
 9 & 45 \\
 \hdashline
 10 & 42 \\
 \hdashline
\end{array}"

"P(37\leq X\leq 50)=\dfrac{9}{10}=0.9"